# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ 2.2 What is Information Theoretic (MDL) probing?
3 Methodology
3.1 Models
For the majority of our experiments, we compare our 25 MultiContriever models to the 25 MultiBerts models (Sellam et al., 2022). We access the MultiBerts via huggingface333e.g. https://huggingface.co/google/MultiBerts-seed_[SEED] and train the contrievers via modifying the repository released in Izacard et al. (2022). We use the same contrastive training data as Izacard et al. (2022), to maximise comparability. This comprises a 50/50 mix of Wikipedia and CCNet from 2019. As a result, five of the fourteen performance datasets involve temporal generalisation, since they postdate both the MultiContriever and the MultiBert training data. This most obviously affects the TREC-COVID dataset (QA), though also four additional datasets: Touché-2020 (argumentation), SCIDOCS (citation prediction), and Climate-FEVER and Scifact (fact-checking). Further details on contriever training and infrastructure are in Appendix A.
3.2 Probing Datasets
We verify that results are not dataset specific, or the result of dataset artifacts, by using two probing datasets. First the BiasinBios dataset (De-Arteaga et al., 2019), which contains biographies from the web annotated with labels of the subject’s binary gender (male, female) and biography topic (lawyer, journalist, etc). We also use the Wikipedia dataset from md_gender (Dinan et al., 2020), which contains Wikipedia pages about people, annotated with binary gender labels. [^4] This dataset does contain non-binary labels, but there are few (0.003%, or 1̃80 examples out of 6 million). Uniform codelength ($dataset\_size*log2(num\_classes)$) affects information theoretic probing; additional class with very few examples can significantly affect results. This dataset was also noisier, making small data subsets less trustworthy. For gender labels, BiasinBios is close to balanced, with 55% male and 45% female labels, but Wikipedia is very imbalanced, with 85% male and 15% female. For topic labels, BiasinBios has a long-tail zipfian distribution over 28 professions, with professor and physician together as a third of examples and rapper and personal trainer as 0.7%. Examples from both datasets can be found in Appendix B.
3.3 Evaluation Datasets and Metrics
We evaluate on the BEIR benchmark, which covers retrieval for seven different tasks (fact-checking, citation prediction, duplicate question retrieval, argument retrieval, question answering, bio-medical information retrieval, and entity retrieval). [^6] The BEIR benchmark itself contains two additional tasks, tweet retrieval, and news retrieval, but these datasets are not publicly available.. We initially analysed all standard metrics used in BEIR and TREC (e.g. NDCG, Recall, MAP, MRR, @10 and @100). We observed similar trends across all metrics, somewhat to our surprise, since many retrieval papers focus on the superiority of a particular metric (Wang et al., 2013). We thus predominantly report NDCG@10, but more metrics (NDCG@100, and Recall@100) are included in Appendix G.
For allocational fairness evaluation, we create NQ-gender, a subset of Natural Questions (NQ) about entities, annotated with male, female, and neutral (no gender). Further details on annotation in Appendix C. We measure allocational fairness as the difference between the female and male query performance. We use the neutral/no gender entity queries as a control to make sure the system performs normally on this type of query.
@@ 3.3 Evaluation Datasets and Metrics
4 Results
4.1 Q1: Information Extractability
@@ 4.2 Q2: Sensitivity to Random Initialisation
We analysed the distribution of performance by dataset for 24 seeds, as both generalisation and fairness are sensitive to initialisation in MLMs (Sellam et al., 2022). [^7] Seed 13 (ominously) is excluded from our analysis because of extreme outlier behaviour, which was not reported in (Sellam et al., 2022). We investigated this behaviour, and it is fascinating, but orthogonal to this work, so we have excluded the seed from all analysis. Our investigation can be found in Appendix E and should be of interest to researchers investigating properties of good representations (e.g. anisotropy) and of random initialisations. Figure 4 shows this data, broken out by dataset, with a dashed line at previous reference performance (Izacard et al., 2022).
A few things are notable: first, there is a large range of benchmark performance across seeds with for identical contrastive losses. During training, MultiContrievers converge to the same accuracy (Appendix A) and (usually) have the same aggregate BEIR performance reported in Izacard et al. (2022). However, the range of scores per dataset is often quite large, and for some datasets the original reference Contriever is in the tail of the distribution: e.g in Climate-Fever (row 1 column 2) it performs much worse than all 24 models. It is also worse than almost all models for Fiqa and Arguana. [^8] For Fiqa 19 models are up to 2.5 points better, for Arguana 20 models are up to 6.3 points better. Nothing changed between the different MultiContrievers except the random seed for MultiBert initialisation, and the random seed for the data shuffle for contrastive fine-tuning. [^9] There are a few small differences between the original released BERT, which Contriever was trained on, and the MultiBerts, which we trained on, detailed in Sellam et al. (2022). But not between our 25 MultiBerts.
@@
Third, the best and worst model across the BEIR benchmark datasets is not consistent (Figure 5); not only is the range large across seeds but the ranking of each seed is very variable. The best model on average, seed 24, is top-ranked on only one dataset, and the second-best average model, seed 2, is best on no individual datasets. The best or worst model on any given dataset is almost always the best or worst on only that dataset and none of the other 14. Sometimes, the best model on one dataset is worst on another, e.g. seed 4 is best on NQ and worst on FiQA, seed 5 is best on Scifact and worst on Scidocs. [^11] This best-worst flip exists for seeds 8, 18, and 23 also.. Even seed 10, which is the only model that is worst on more than 2 datasets (it is worst on 6) is still best on TREC-Covid. [^12] This is to be taken with a grain of salt - that dataset is interesting for generalisation (as these models are trained on only pre-Covid data), but it is only 50 datapoints. We note also that analysis on seed 13 revealed that seed 10 was also unusual, that analysis can also be found in E.
@@
As an addendum, we note that Sellam et al. (2022) did extensive experiments with both random initialisation and data shuffle, and found initialisation to matter more. We did our own experiments to this effect where we trained five MultiContrievers from the same MultiBert initialisation with different data shuffles, from the best, worst, and middle performing seeds. This additional analysis is in Appendix D.
@@ 4.3 Q3: Correlation between Extractability & Performance
We tested for correlations across all datasets and common metrics, and present a selection here (Fig 3). Neither NQ (Fig 2(a)) nor MSMARCO (Fig 2(b)) correlate with compression metrics. NQ and MSMARCO are the most widely used of the BEIR benchmark datasets, and we hypothesised them to be most likely to correlate. Both are search engine queries (from Google and Bing, respectively) and contain queries that require occupation-type information (what is cabaret music?, MSMARCO) and that require gender information (who is the first foreign born first lady?, NQ). However, as the dispersed points on the scatterplots show (Figures 2(a), 2(b)), neither piece of information correlates to performance on either dataset. NQ and MSMARCO are representative; we include plots for all datasets in Appendix F.
This result was somewhat surprising; since the contriever training both regularises and increases extractability of gender and occupation, we might expect this to be important for the task. But perhaps it is relevant for only the contrastive objective, and not for the retrieval benchmark. Alternatively, it is possible that this information is important, but only up to some threshold that MultiContriever models exceed. Finally, it’s possible that this information doesn’t matter for most queries in these datasets, and so there is some correlation but it is lost, as these datasets are extremely large. This is somewhat supported by the exception cases with correlations being smaller, more curated datasets (F), and so we investigated this as the most tractable to implement.
Figure 6: INLP experiments
@@ 6 Limitations
Appendix A Contriever Training
Each MultiContriever was trained across 4 nodes with 8 GPUs per node (32 GPUs total) for on average 2.5 days. Each MultiContriever was trained for the full 500,000 steps, and checkpointed often; but in all but one seed the best performing checkpoint was the final one (so for that one we use the model at 450,000 steps). This is excepting seed 13, which was anomalous in many other ways (see E).
@@ Appendix A Contriever Training
Appendix B Probing Datasets
Appendix C Annotation of NQ gender subset
Appendix D Data Shuffle Experiments
Appendix E Seed 13
(a) MultiBerts (mean)
(b) MultiBerts (var)
(c) MultiContriever (mean)
(d) MultiContriever (var)
(e) MultiBerts (mean)
(f) MultiBerts (var)
(g) MultiContriever (mean)
(h) MultiContriever (var)
Appendix F Full set of results for correlation between extractability and performance
Appendix G Additional metrics

### v2

# MultiContrievers: Analysis of Dense Retrieval Representations

@@ 2.2 What is Information Theoretic (MDL) probing?
3 Methodology
3.1 Models
For the majority of our experiments, we compare our 25 MultiContriever models to the 25 MultiBerts models (Sellam et al., 2022). We access the MultiBerts via huggingface333e.g. https://huggingface.co/google/MultiBerts-seed_[SEED] and train the contrievers via modifying the repository released in Izacard et al. (2022). We use the same contrastive training data as Izacard et al. (2022), to maximise comparability. This comprises a 50/50 mix of Wikipedia and CCNet from 2019. As a result, five of the fourteen performance datasets involve temporal generalisation, since they postdate both the MultiContriever and the MultiBert training data. This most obviously affects the TREC-COVID dataset (QA), though also four additional datasets: Touché-2020 (argumentation), SCIDOCS (citation prediction), and Climate-FEVER and Scifact (fact-checking). Further details on contriever training and infrastructure are in Appendix G.
3.2 Probing Datasets
We verify that results are not dataset specific, or the result of dataset artifacts, by using two probing datasets. First the BiasinBios dataset (De-Arteaga et al., 2019), which contains biographies from the web annotated with labels of the subject’s binary gender (male, female) and biography topic (lawyer, journalist, etc). We also use the Wikipedia dataset from md_gender (Dinan et al., 2020), which contains Wikipedia pages about people, annotated with binary gender labels. [^4] This dataset does contain non-binary labels, but there are few (0.003%, or 1̃80 examples out of 6 million). Uniform codelength ($dataset\_size*log2(num\_classes)$) affects information theoretic probing; additional class with very few examples can significantly affect results. This dataset was also noisier, making small data subsets less trustworthy. For gender labels, BiasinBios is close to balanced, with 55% male and 45% female labels, but Wikipedia is very imbalanced, with 85% male and 15% female. For topic labels, BiasinBios has a long-tail zipfian distribution over 28 professions, with professor and physician together as a third of examples and rapper and personal trainer as 0.7%. Examples from both datasets can be found in Appendix A.
3.3 Evaluation Datasets and Metrics
We evaluate on the BEIR benchmark, which covers retrieval for seven different tasks (fact-checking, citation prediction, duplicate question retrieval, argument retrieval, question answering, bio-medical information retrieval, and entity retrieval). [^6] The BEIR benchmark itself contains two additional tasks, tweet retrieval, and news retrieval, but these datasets are not publicly available.. We initially analysed all standard metrics used in BEIR and TREC (e.g. NDCG, Recall, MAP, MRR, @10 and @100). We observed similar trends across all metrics, somewhat to our surprise, since many retrieval papers focus on the superiority of a particular metric (Wang et al., 2013). We thus predominantly report NDCG@10, but more metrics (NDCG@100, and Recall@100) are included in Appendix F.
For allocational fairness evaluation, we create NQ-gender, a subset of Natural Questions (NQ) about entities, annotated with male, female, and neutral (no gender). Further details on annotation in Appendix B. We measure allocational fairness as the difference between the female and male query performance. We use the neutral/no gender entity queries as a control to make sure the system performs normally on this type of query.
@@ 4.2 Q2: Sensitivity to Random Initialisation
We analysed the distribution of performance by dataset for 24 seeds, as both generalisation and fairness are sensitive to initialisation in MLMs (Sellam et al., 2022). [^7] Seed 13 (ominously) is excluded from our analysis because of extreme outlier behaviour, which was not reported in (Sellam et al., 2022). We investigated this behaviour, and it is fascinating, but orthogonal to this work, so we have excluded the seed from all analysis. Our investigation can be found in Appendix D and should be of interest to researchers investigating properties of good representations (e.g. anisotropy) and of random initialisations. Figure 5 shows this data, broken out by dataset, with a dashed line at previous reference performance (Izacard et al., 2022).
A few things are notable: first, there is a large range of benchmark performance across seeds with for identical contrastive losses. During training, MultiContrievers converge to the same accuracy (Appendix G) and (usually) have the same aggregate BEIR performance reported in Izacard et al. (2022). However, the range of scores per dataset is often quite large, and for some datasets the original reference Contriever is in the tail of the distribution: e.g in Climate-Fever (row 1 column 2) it performs much worse than all 24 models. It is also worse than almost all models for Fiqa and Arguana. [^8] For Fiqa 19 models are up to 2.5 points better, for Arguana 20 models are up to 6.3 points better. Nothing changed between the different MultiContrievers except the random seed for MultiBert initialisation, and the random seed for the data shuffle for contrastive fine-tuning. [^9] There are a few small differences between the original released BERT, which Contriever was trained on, and the MultiBerts, which we trained on, detailed in Sellam et al. (2022). But not between our 25 MultiBerts.
@@
Third, the best and worst model across the BEIR benchmark datasets is not consistent (Figure 4); not only is the range large across seeds but the ranking of each seed is very variable. The best model on average, seed 24, is top-ranked on only one dataset, and the second-best average model, seed 2, is best on no individual datasets. The best or worst model on any given dataset is almost always the best or worst on only that dataset and none of the other 14. Sometimes, the best model on one dataset is worst on another, e.g. seed 4 is best on NQ and worst on FiQA, seed 5 is best on Scifact and worst on Scidocs. [^11] This best-worst flip exists for seeds 8, 18, and 23 also.. Even seed 10, which is the only model that is worst on more than 2 datasets (it is worst on 6) is still best on TREC-Covid. [^12] This is to be taken with a grain of salt - that dataset is interesting for generalisation (as these models are trained on only pre-Covid data), but it is only 50 datapoints. We note also that analysis on seed 13 revealed that seed 10 was also unusual, that analysis can also be found in D.
@@
As an addendum, we note that Sellam et al. (2022) did extensive experiments with both random initialisation and data shuffle, and found initialisation to matter more. We did our own experiments to this effect where we trained five MultiContrievers from the same MultiBert initialisation with different data shuffles, from the best, worst, and middle performing seeds. This additional analysis is in Appendix C.
@@ 4.3 Q3: Correlation between Extractability & Performance
We tested for correlations across all datasets and common metrics, and present a selection here (Fig 3). Neither NQ (Fig 3(a)) nor MSMARCO (Fig 3(b)) correlate with compression metrics. NQ and MSMARCO are the most widely used of the BEIR benchmark datasets, and we hypothesised them to be most likely to correlate. Both are search engine queries (from Google and Bing, respectively) and contain queries that require occupation-type information (what is cabaret music?, MSMARCO) and that require gender information (who is the first foreign born first lady?, NQ). However, as the dispersed points on the scatterplots show (Figures 3(a), 3(b)), neither piece of information correlates to performance on either dataset. NQ and MSMARCO are representative; we include plots for all datasets in Appendix E.
This result was somewhat surprising; since the contriever training both regularises and increases extractability of gender and occupation, we might expect this to be important for the task. But perhaps it is relevant for only the contrastive objective, and not for the retrieval benchmark. Alternatively, it is possible that this information is important, but only up to some threshold that MultiContriever models exceed. Finally, it’s possible that this information doesn’t matter for most queries in these datasets, and so there is some correlation but it is lost, as these datasets are extremely large. This is somewhat supported by the exception cases with correlations being smaller, more curated datasets (E), and so we investigated this as the most tractable to implement.
@@ 6 Limitations
Appendix A Probing Datasets
Appendix B Annotation of NQ gender subset
Appendix C Data Shuffle Experiments
Appendix D Seed 13
(a) MultiBerts (mean)
(b) MultiBerts (var)
(c) MultiContriever (mean)
(d) MultiContriever (var)
(e) MultiBerts (mean)
(f) MultiBerts (var)
(g) MultiContriever (mean)
(h) MultiContriever (var)
Appendix E Full set of results for correlation between extractability and performance
Appendix F Additional metrics
@@
Appendix G Contriever Training
Each MultiContriever was trained across 4 nodes with 8 GPUs per node (32 GPUs total) for on average 2.5 days. Each MultiContriever was trained for the full 500,000 steps, and checkpointed often; but in all but one seed the best performing checkpoint was the final one (so for that one we use the model at 450,000 steps). This is excepting seed 13, which was anomalous in many other ways (see D).